How do I solve this problem
Eric's gas tank is 1/2 full. After he buys 2 gallons of gas, it is 2/3 full. How many gallons can Eric's tank hold?

Answer by rfadrogane(214)

Eric's gas tank is 1/2 full. After he buys 2 gallons of gas, it is 2/3 full. How many gallons can Eric's tank hold?
Sol'n:
this problem is a ratio and proportion:
2/3-1/2=1/6 --this corresponds to the 2 gallons.
let x - the total volume of the tank
1/x=(1/6)/2
x=12 gallons ----answer
^_^